OpenStudy (anonymous):

I need help finding the LCM and the GCF for the ages 10, 31 and 44

OpenStudy (amistre64):

\[\frac{1}{10}+\frac{1}{31}+\frac{1}{44}=N\] \[\frac{1(10)}{10}+\frac{1(10)}{31}+\frac{1(10)}{44}=(10)N\] \[1+\frac{10}{31}+\frac{5}{22}=10N\] \[1(22)+\frac{10(22)}{31}+\frac{5(22)}{22}=10(22)N\] \[22+\frac{220}{31}+5=220N\] the LCM looks to be 220(31) maybe

myininaya (myininaya):

10,31,44 have no common factors 10=2*5 31=31 44=2*2*11 lcm(10,31,44)=2*2*5*11*31=2860

myininaya (myininaya):

what is N?

OpenStudy (zarkon):

lcm 6820

myininaya (myininaya):

because i can't multiply lol

OpenStudy (zarkon):

;)

OpenStudy (amistre64):

N is just some number

OpenStudy (amistre64):

I use it as a place holder :) so I can algebra it

myininaya (myininaya):

ok amistre yes 2*2*5*11*31=6820

OpenStudy (amistre64):

yay!!

myininaya (myininaya):

when you put this in the calculator make sure you don't put in 13 instead of 31 i swear i looked at those numbers like 5 times and everything looked okay and then finally my dyslexia seen it was 31 not 13

myininaya (myininaya):

i was like my calculator must have gave me the wrong answer i thought it had dyslexia

OpenStudy (amistre64):

had a bit of sulfur in your eyes ;)

myininaya (myininaya):

i don't have my glasses on right now

myininaya (myininaya):

i'm blow drying my hair and fixing to flat iron it :)

OpenStudy (amistre64):

multitasking and math, its a bad combonation

OpenStudy (amistre64):

math and typing is hard enough ;)

OpenStudy (anonymous):

i will attach what i have from my instructor